### v2

@@ 6 Concluding remarks
Applications of the above theorems to concrete interacting QFT models were provided, over continuum flat Euclidean signature spacetimes, in dimensions 1 to 4 (Section 5). In particular, we prove that in arbitrary dimensional spacetimes, an interacting QFT model defined by a free Klein–Gordon kinetic term + an interaction potential which is bounded both from below and above, is nonperturbatively renormalizable (Theorem 28, Corollary 29). Such models are defined e.g. by the “Weidling” potential $V(\varphi)=g\,\int\frac{1}{1+l^{4}\,\varphi^{4}}\,\varphi^{4}$, or by a sine–Gordon model. Possible triviality of these models is still an open question (for very recent advances on that question, see [23]). The nonperturbative renormalizability of further relevant models, such as the competing Higgs model defined by the potential $V(\varphi,\psi)=g\,\int(\varphi^{2}-\psi^{2})^{2}$ also could not be decided, but there are indications for its renormalizability (it avoids the no-go Theorem 32, which obstructs high dimensional $\varphi^{4}$ models). The results proved in the paper were also discussed in the light of known examples, in particular the $\varphi^{4}$ models over 1 to 4 dimensional spacetimes [38, 21, 22, 39].
[…]
Appendix
@@ Appendix A On the supports of Gaussian measures
For a sigma-additive non-negative valued measure $\gamma$ on the Borel sets of some topological space there is the well known notion of topological support, denoted by $\mathop{\mathrm{supp}}\nolimits(\gamma)$, see e.g. [27] Definition A.3.14. It is defined as the intersection of those closed sets whose complements are $\gamma$-measure zero. By construction, the topological support of a measure is a closed set, and is a generalization of the support notion of the continuous functions on topological spaces. Not surprisingly, typical Gaussian measures turning up in QFT are so-called non-degenerate, meaning that their topological support is the full space $\mathcal{S}^{\prime}$. Since, however, in QFT the considered density functions $\mathrm{e}^{-V}$ to be integrated against the reference Gaussian measure $\gamma$ are not continuous, only measurable, the topological support of the reference measure $\gamma$ is not very informative on the non-vanishing of the overlap integral of $\mathrm{e}^{-V}$ and $\gamma$.
Due to the above complication, another notion of support is also used in the literature. A Borel set $B$ is said to support $\gamma$ if its complement has zero $\gamma$-measure. Clearly, if one can find some Borel set $B$ such that its $\gamma$-measure is nonzero, and the density function $\mathrm{e}^{-V}$ is nonwhere zero on $B$, then their product $\mathrm{e}^{-V}\cdot\gamma$ is not the zero measure, i.e. it defines a proper probability measure modulo normalization, meaning that $V$ describes a renormalizable QFT model.
@@
In the following we recall some theorems on Borel sets from $\mathcal{S}^{\prime}(\mathbb{R}^{N})$ which are known to support a Gaussian measure $\gamma$ with Fourier transform $j\mapsto\mathrm{e}^{-\frac{1}{2}(j\,|\,\mathtt{K}\star j)}$ here $\mathtt{K}$ denoting the maximally symmetric fundamental solution of a Euclidean Klein–Gordon operator $(-\Delta+m^{2})$ on $\mathbb{R}^{N}$. Also, theorems are recalled on those Borel sets which are known to have zero $\gamma$-measure against such Gaussian. See an excellent review in [2, 40, 27].
@@
    Two Gaussian measures on the same locally convex topological vector space which are obtained from the same Gaussian measure via pushforward by field rescaling factors $z,z^{\prime}\in\mathbb{R}^{+}$, are not mutually singular precisely when $z^{\prime}=z$ (see e.g. [27] Example 2.7.4).

### v3

@@ 6 Concluding remarks
Applications of the above theorems to concrete interacting QFT models were provided, over continuum flat Euclidean signature spacetimes, in dimensions 1 to 4 (Section 5). In particular, we prove that in arbitrary dimensional spacetimes, an interacting QFT model defined by a free Klein–Gordon kinetic term + an interaction potential which is bounded both from below and above, is nonperturbatively renormalizable (Theorem 28, Corollary 29). Such models are defined e.g. by the “basin-shaped” potential $V(\varphi)=g\,\int\frac{1}{1+l^{4}\,\varphi^{4}}\,\varphi^{4}$, or by a sine–Gordon model. (On the triviality question of some of these models, see the very recent results of [27].) The nonperturbative renormalizability of further relevant models, such as the competing Higgs model defined by the potential $V(\varphi,\psi)=g\,\int(\varphi^{2}-\psi^{2})^{2}$ also could not be decided, but there are indications for its renormalizability (it avoids the no-go Theorem 32, which obstructs high dimensional $\varphi^{4}$ models). The results proved in the paper were also discussed in the light of known examples, in particular the $\varphi^{4}$ models over 1 to 4 dimensional spacetimes [17, 18, 19, 20].
Acknowledgements
The authors would like to thank to János Balog, Gergely Fejős, Antal Jakovác, Bálint Tóth, Jan Dereziński and Markus Fröb for discussions and valuable insights.
This work was supported in part by the Hungarian Scientific Research fund, NKFI K-138136-138152, K-142423, ADVANCED-150038, ADVANCED-150059. Z. Tarcsay was supported by Project no. TKP2021-NVA-09, which has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme.
@@ Appendix A On the supports of Gaussian measures
For a sigma-additive non-negative valued measure $\gamma$ on the Borel sets of some topological space there is the well known notion of topological support, denoted by $\mathop{\mathrm{supp}}\nolimits(\gamma)$, see e.g. [31] Definition A.3.14. It is defined as the intersection of those closed sets whose complements are $\gamma$-measure zero. By construction, the topological support of a measure is a closed set, and is a generalization of the support notion of the continuous functions on topological spaces. Not surprisingly, typical Gaussian measures turning up in QFT are so-called non-degenerate, meaning that their topological support is the full space $\mathcal{S}^{\prime}$. Since, however, in QFT the considered density functions $\rme^{-V}$ to be integrated against the reference Gaussian measure $\gamma$ are not continuous, only measurable, the topological support of the reference measure $\gamma$ is not very informative on the non-vanishing of the overlap integral of $\rme^{-V}$ and $\gamma$.
Due to the above complication, another notion of support is also used in the literature. A Borel set $B$ is said to support $\gamma$ if its complement has zero $\gamma$-measure. Clearly, if one can find some Borel set $B$ such that its $\gamma$-measure is nonzero, and the density function $\rme^{-V}$ is nowhere zero on $B$, then their product $\rme^{-V}\cdot\gamma$ is not the zero measure, i.e. it defines a proper probability measure modulo normalization, meaning that $V$ describes a renormalizable QFT model.
@@
In the following we recall some theorems on Borel sets from $\mathcal{S}^{\prime}(\mathbb{R}^{N})$ which are known to support a Gaussian measure $\gamma$ with Fourier transform $j\mapsto\rme^{-\frac{1}{2}(j\,|\,\mathtt{K}\star j)}$ here $\mathtt{K}$ denoting the maximally symmetric fundamental solution of a Euclidean Klein–Gordon operator $(-\Delta+m^{2})$ on $\mathbb{R}^{N}$. Also, theorems are recalled on those Borel sets which are known to have zero $\gamma$-measure against such Gaussian. See an excellent review in [2, 43, 31].
@@
    Two Gaussian measures on a locally convex topological vector space which are obtained from the same Gaussian measure (having non-finite dimensional support) via pushforward by field rescaling factors $z,z^{\prime}\in\mathbb{R}^{+}$, are not mutually singular precisely when $z^{\prime}=z$ (see e.g. [31] Example 2.7.4).